# Shadow and Polarization Images of Rotating Black Holes in Kalb-Ramond Gravity Illuminated by Several Thick Accretion Disks

### v1

@@ Abstract
Using ray-tracing techniques, this paper investigates the optical and polarization images of rotating black holes in Kalb-Ramond (KR) gravity illuminated by thick accretion disks. We consider two accretion disk models: the phenomenological model and the HOU disk model, where the phenomenological model includes both isotropic and anisotropic radiation. In all models, the external bright rings corresponding to the high-order image and the internal dark region associated with the event horizon are observed. At high observational inclinations, the inner shadows are obscured by the radiation from the equatorial plane, which is significantly different from the thin accretion disk model. The primary distinction between isotropic and anisotropic radiation is that the latter causes distortion of the high-order image in the vertical direction, resulting in an elliptical structure. For the HOU disk model, due to certain regions are geometrically thinner under the conical approximation, the high-order images are narrower compared to the phenomenological model. Furthermore, we find that an increase in the rotational parameter $a$ leads to an asymmetry in the intensity distribution of the high-order image, while an increase in the spontaneous Lorentz violating parameters, $\varsigma$ and $\varpi$, results in a decrease in the size of the high-order image. In the polarization image, the linear polarization is found to be significantly influenced by the intensity, while it is relatively less affected by the parameters $\varsigma$ and $\varpi$. The electric vector position angle is mainly affected by the parameters $\varsigma$ and $\varpi$.
@@ 1 Introduction
Supermassive black holes accrete hot, magnetized plasma, forming a bright accretion disk. The frequency observed in the image is dominated by thermal synchrotron radiation [8, 9]. Meanwhile, for high-spin black holes, electromagnetic energy may trigger relativistic jets [10], known as ”funnel wall jets” (FWJ). Theoretically, both the accretion disk and the jets can act as light sources, generating black hole shadow images. The study of black hole shadows has a long history [11, 12], including spherically symmetric accretion models [13, 14, 15], optical and geometric thin accretion disk models [16, 17, 18, 19], and holographic Einstein rings [20, 21, 22]. Based on EHT observational data, many studies have also attempted to constrain the parameters of black hole models [23, 24, 25, 26, 27]. Current studies on black hole shadows mostly assume the accretion disk to be optically and geometrically thin. However, EHT and related observational studies suggest that in strong gravitational fields, the accretion flow near supermassive black holes, due to suppressed vertical cooling and matter compression, may become geometrically thick and optically thin [3, 9, 4, 28, 29]. In this case, parameters such as electron number density, electron temperature, and magnetic field structure must be considered. In theoretical studies of geometrically thick accretion disks, most work uses the radiatively inefficient accretion flow (RIAF) model with relatively low radiation efficiency, where the vertically-averaged electron number density and temperature typically follow a power-law distribution with radius [30, 31, 32, 33, 34].
@@
Advances in experimental techniques have not only enabled us to verify the applicability of General Relativity (GR) with higher precision and resolution [45, 46], but have also driven the process of modifying GR and exploring alternative gravitational theories. An important method for modifying the GR is to alter the Einstein action. The introduction of the Kalb-Ramond (KR) field is a key approach [47]. The KR field, as a two-form quantum field, is believed to be related to the closed string excitation in heterotic string theory [48]. Studies in the KR field have led to many important conclusions, such as the derivation of a third-order antisymmetric tensor, which is interpreted as the source of spacetime torsion and the intrinsic angular momentum of celestial bodies [49, 50]. Further research on the gravitational effects of the KR field and its influence on particles can be found in [51, 52, 53, 54].
@@
The structure of this paper is organized as follows. Section 2 briefly introduces rotating black holes in KR gravity and presents the null geodesic equation near the black hole. Sections 3 and 4 introduce the electron radiation model and the accretion disk model, respectively, where the electron radiation model includes both isotropic and anisotropic radiation, and the accretion disk model includes the phenomenological model and the HOU disk model. Section 5 presents the basic theoretical background for polarization imaging in the case of anisotropic radiation. Section 6 provides the numerical results. Finally, Section 7 concludes with a summary and discussion. Unless otherwise specified, we adopt geometric units, i.e., $c=G=1$, where $c$ is the vacuum speed of light and $G$ is the gravitational constant.
@@ 2 Review of Black Holes in Kalb-Ramond Gravity
$\omega=\frac{d\phi}{dt}=-\frac{g_{t\phi}}{g_{\phi\phi}}=\frac{2Mar-\frac{\varsigma a}{r^{-2(\varpi-1)/\varpi}}}{[(r^{2}+a^{2})^{2}-a^{2}\Delta]}.$ (6)
@@ 3.2 Synchrotron Radiation
$e_{(k)}^{\mu}=-\left(\frac{k^{\mu}}{u^{\nu}k}+u^{\mu}\right),\quad e_{(b)}^{\mu}=\frac{b^{\mu}}{b},$ (29)
@@ 4 Accretion Disk Models
For the non-polarized case, two geometrically thick and optically thin accretion disk models can be considered: the phenomenological model [62] and the HOU disk model [63, 64]. The former is highly consistent with general relativistic magnetohydrodynamic (GRMHD) simulations and has successfully reproduced the overall morphological features of M87* [9]. However, it neglects outflows, non-thermal particles, and the full GRMHD dynamics, thus limiting its application in polarization studies. The latter assumes that the fluid near the event horizon is primarily accelerated by gravity and provides explicit expressions for the thermodynamic variables and magnetic field configuration, offering a better description of the morphology and dynamics of the geometrically thick accretion flow in the near-horizon region of the black hole.
4.1 Phenomenological Model
@@
where the dimensionless quantity $\rho=n_{e}(m_{p}c^{2})$ represents the fluid mass density. For the accretion disk model considered in this paper, the order of magnitude of $\sigma$ is $\sigma\sim 0.1$ [65]. For the phenomenological model, both isotropic radiation (35) and anisotropic radiation (33) can be considered.
@@
4.2 HOU Disk Model
The HOU disk model is a steady-state, axisymmetric accretion model [64, 63]. In this model, the fluid is treated as electrically neutral and the plasma is assumed to be fully ionized into electrons and protons. The accreting particle is confined to constant $\theta$ surface, i.e., $u^{\theta}\equiv 0$. In this case, the mass conservation equation is
@@ 4.2 HOU Disk Model
$\Omega_{b}$ is the angular velocity of the magnetic field, and in this paper, we take $\Omega_{b}=0.3\Omega_{h}$, where $\Omega_{h}={a}/(2r_{h})$ is the angular velocity of the rotating black hole. For the HOU disk model, only anisotropic radiation (33) is considered. For the fluid’s four-velocity, we can use the ballistic approximation, where the fluid moves along the geodesics [64], or the infalling motion mentioned in (47). We will choose the later in this paper.
@@ 6 Numerical Results
6.1 Phenomenological Model under Isotropic Radiation
We first discuss the impact of the spin parameter $a$ on the phenomenological model under isotropic radiation. We set the spontaneous Lorentz violating parameters $\varpi=\varsigma=0.1$, the observer inclination angle is $\theta_{o}=65^{\circ}$. The numerical results are shown in Figure 1. We observe a bright ring structure in all the images, corresponding to higher-order images, where photons orbit the black hole one or more times before reaching the observer. These higher-order images are a result of the strong gravitational lensing effect, where the black hole’s gravity bends the light path, allowing some photons to bypass the black hole and reach the observer. Beyond this ring structure, the region with nonzero intensity corresponds to the primary image, where photons travel directly from the accretion disk to the observer. Notably, for all parameters, there is a region of decreasing intensity inside the higher-order images, which originates from the black hole’s event horizon. In the case of a geometrically thin accretion disk, the particle is confined to the equatorial plane, so the event horizon appears as a clearly defined black region in the image (the ”inner shadow”), which may be captured by the EHT [74]. However, for a geometrically thick accretion disk, this region may be obscured by radiation from outside the equatorial plane, making it difficult to distinguish. Compared to thin disks, thick disks are more physically realistic, suggesting that direct imaging of the black hole’s event horizon remains challenging.
@@ 6.1 Phenomenological Model under Isotropic Radiation
Figure 1: The impact of the spin parameter $a$ on the phenomenological model under isotropic radiation. Fixed parameters are $\varpi=\varsigma=0.1$, and $\theta_{o}=65^{\circ}$.
@@
Figure 3: Black hole shadow images under the phenomenological model with isotropic radiation. The parameters are fixed at $a=0.1$, $\theta_{o}=65^{\circ}$.
@@
Figure 4: Intensity distributions along the x-axis and y-axis under the phenomenological model with isotropic radiation. The red, green, blue, and orange colors correspond to $\varsigma=0.69,0.79,0.89,0.99$, respectively. The parameters are fixed at $a=0.1$, $\varpi=0.9$, $\theta=65^{\circ}$.
@@
Figure 5: Intensity distributions along the x-axis and y-axis under the phenomenological model with isotropic radiation. The red, green, and blue colors correspond to $\varpi=0.3,0.6,0.9$, respectively. The parameters are fixed at $a=0.1$, $\varsigma=0.99$, $\theta_{o}=65^{\circ}$.
6.2 Phenomenological Model under Anisotropic Radiation
Figures 6 and 7 show the impact of the spin parameter $a$ on the phenomenological model under anisotropic radiation, along with the corresponding blurred images. For isotropic radiation, when $a$ is small, the higher-order images are approximately circular, but under anisotropic radiation, they significantly deform into elliptical shapes. The main reason for this difference is the different fitting functions for $\mathcal{F}(x)$. Similar to Figure 1, as $a$ increases, the intensity on the left side of the ellipse increases. Figure 8 shows the impact of $\varpi$ and $\varsigma$ on the shadow image. It can be observed that increasing $\varpi$ and $\varsigma$ reduces the size of the ellipse. Figures 9 and 10 show the corresponding intensity distributions. For both the x-axis and the y-axis, under anisotropic radiation, the distance between the two peaks is smaller than in the case of isotropic radiation. The above analysis indicates that the radiation type affects both the shape and size of the higher-order images.
@@ 6.2 Phenomenological Model under Anisotropic Radiation
Figure 6: The impact of the spin parameter $a$ on the phenomenological model under anisotropic radiation. The fixed parameters are $\varpi=\varsigma=0.1$, and $\theta_{o}=65^{\circ}$.
@@
Figure 8: Black hole shadow images under the phenomenological model with anisotropic radiation. The fixed parameters are $a=0.1$, $\theta_{o}=65^{\circ}$.
@@
Figure 9: Intensity distributions along the x-axis and y-axis under the phenomenological model with anisotropic radiation. The red, green, blue, and orange colors correspond to $\varsigma=0.69,0.79,0.89,0.99$, respectively. The fixed parameters are $a=0.1$, $\varpi=0.9$, $\theta=65^{\circ}$.
@@
Figure 10: Intensity distributions along the x-axis and y-axis under the phenomenological model with anisotropic radiation. The red, green, and blue colors correspond to $\varpi=0.3,0.6,0.9$, respectively. The fixed parameters are $a=0.1$, $\varsigma=0.99$, $\theta_{o}=65^{\circ}$.
6.3 HOU Disk Model
Figure 11 shows the impact of the spin parameter $a$ on the HOU disk model, and Figure 12 presents the corresponding blurred images. When $a$ is small, the higher-order images appear circular, similar to the case of isotropic radiation. However, unlike the isotropic case, the higher-order images in the HOU disk model feature a complete region of intensity reduction inside them. This indicates that, compared to the phenomenological model, the radiation from outside the equatorial plane in the HOU disk model has a weaker effect in obscuring the event horizon. Comparing the phenomenological model and the HOU disk model, we find that the latter is closer to the EHT observational results. As shown in Figure 13, the impact of $\varpi$ and $\varsigma$ on the higher-order images in the HOU disk model is similar to that in the phenomenological model. However, the region corresponding to the primary image is significantly smaller in the HOU disk model. This conclusion is also evident from the intensity distribution diagrams in Figures 14 and 15. For both the x-axis and y-axis, the peaks corresponding to the higher-order images in the HOU disk model are sharper, and the intensity rapidly approaches zero beyond the peaks. This reflects that, in the HOU disk model, the intensity contribution primarily originates from the higher-order images. In contrast, in the phenomenological model, the intensity of the primary image outside the higher-order images remains relatively high.
@@ 6.3 HOU Disk Model
Figure 11: The impact of the spin parameter $a$ on the HOU disk model. The fixed parameters are $\varpi=\varsigma=0.1$, and $\theta_{o}=65^{\circ}$.
@@
Figure 13: Black hole shadow image under the HOU disk model. The parameters are fixed at $a=0.1,\theta_{o}=65^{\circ}$.
@@
Figure 14: Intensity distribution along the x-axis and y-axis for the HOU disk model. The red, green, blue, and orange colors correspond to $\varsigma=0.69,0.79,0.89,0.99$, respectively. The parameters are fixed as $a=0.1,\varpi=0.9,\theta=65^{\circ}$.
@@
Figure 15: Intensity distribution along the x-axis and y-axis for the HOU disk model. The red, green, and blue colors correspond to $\varpi=0.3,0.6,0.9$, respectively. The parameters are fixed as $a=0.1,\varsigma=0.99,\theta_{o}=65^{\circ}$.
6.4 Polarization Images Illuminated by HOU Disk
For the Polarization imaging, we will take the HOU disk as an example. Figure 16 shows the numerical results for the Stokes parameters $\tilde{I}_{o}$, $\tilde{Q}_{o}$, $\tilde{U}_{o}$, and $\tilde{V}_{o}$. Here, $\tilde{I}_{o}$ reflects the intensity distribution, with arrows representing the linear polarization vector $\vec{f}$. The color and direction of the arrows correspond to the degree of linear polarization $\mathcal{P}_{o}$ and the electric vector position angle $\Phi_{\mathrm{EVPA}}$. Since $\vec{f}\perp\vec{b}$, it can be inferred that the magnetic field is roughly radially distributed. The parameters $\tilde{Q}_{o}$ and $\tilde{U}_{o}$ reach their peaks near the higher-order images and decay rapidly away from this region. The $\tilde{V}_{o}$ parameter indicates right-handed circular polarization on both sides of the higher-order images, with left-handed circular polarization in the remaining regions.

### v2

@@ Abstract
Using ray-tracing techniques, this paper investigates the optical and polarization images of rotating black holes in Kalb-Ramond (KR) gravity illuminated by thick accretion disks. We examine two accretion disk models: the phenomenological radiatively inefficient accretion flow (RIAF) model and the analytical ballistic approximation accretion flow (BAAF) model. The RIAF model incorporates both isotropic and anisotropic radiation. In all models, the external bright rings corresponding to the high-order image and the internal dark region associated with the event horizon are observed. At high observational inclinations, the inner shadows are obscured by the radiation from the equatorial plane, which is significantly different from the thin accretion disk model. The primary distinction between isotropic and anisotropic radiation is that the latter causes distortion of the high-order image in the vertical direction, resulting in an elliptical structure. For the BAAF model, due to certain regions are geometrically thinner under the conical approximation, the high-order images are narrower compared to the RIAF model. Furthermore, we find that an increase in the rotational parameter $a$ leads to an asymmetry in the intensity distribution of the high-order image, while an increase in the spontaneous Lorentz violating parameters, $\varsigma$ and $\varpi$, results in a decrease in the size of the high-order image. In the polarization image, the linear polarization is found to be significantly influenced by the intensity, while it is relatively less affected by the parameters $\varsigma$ and $\varpi$. The electric vector position angle is mainly affected by the parameters $\varsigma$ and $\varpi$.
@@ 1 Introduction
Supermassive black holes accrete hot, magnetized plasma, forming a bright accretion disk. The frequency observed in the image is dominated by thermal synchrotron radiation [8, 9]. Meanwhile, for high-spin black holes, electromagnetic energy may trigger relativistic jets [10], known as ”funnel wall jets” (FWJ). Theoretically, both the accretion disk and the jets can act as light sources, generating black hole shadow images. The study of black hole shadows has a long history [11, 12, 13, 14], including spherically symmetric accretion models [15, 16, 17], optical and geometric thin accretion disk models [18, 19, 20, 21], and holographic Einstein rings [22, 23, 24]. Based on EHT observational data, many studies have also attempted to constrain the parameters of black hole models [25, 26, 27, 28, 29]. Current studies on black hole shadows mostly assume the accretion disk to be optically and geometrically thin. However, EHT and related observational studies suggest that in strong gravitational fields, the accretion flow near supermassive black holes, due to suppressed vertical cooling and matter compression, may become geometrically thick and optically thin [3, 9, 4, 30, 31]. In this case, parameters such as electron number density, electron temperature, and magnetic field structure must be considered. In theoretical studies of geometrically thick accretion disks, most work uses the radiatively inefficient accretion flow (RIAF) model with relatively low radiation efficiency, where the vertically-averaged electron number density and temperature typically follow a power-law distribution with radius [32, 33, 34, 35, 36].
@@
Advances in experimental techniques have not only enabled us to verify the applicability of General Relativity (GR) with higher precision and resolution [47, 48], but have also driven the process of modifying GR and exploring alternative gravitational theories. An important method for modifying the GR is to alter the Einstein action. The introduction of the Kalb-Ramond (KR) field is a key approach [49]. The KR field, as a two-form quantum field, is believed to be related to the closed string excitation in heterotic string theory [50]. Studies in the KR field have led to many important conclusions, such as the derivation of a third-order antisymmetric tensor, which is interpreted as the source of spacetime torsion and the intrinsic angular momentum of celestial bodies [51, 52]. Further research on the gravitational effects of the KR field and its influence on particles can be found in [53, 54, 55, 56, 57, 58, 59].
@@
The structure of this paper is organized as follows. Section 2 briefly introduces rotating black holes in KR gravity and presents the null geodesic equation near the black hole. Sections 3 and 4 introduce the electron radiation model and the accretion disk model, respectively. The electron radiation model incorporates both isotropic and anisotropic radiation, while the accretion disk model includes the phenomenological radiatively inefficient accretion flow (RIAF) model and the analytical ballistic approximation accretion flow (BAAF) model. Section 5 presents the basic theoretical background for polarization imaging in the case of anisotropic radiation. Section 6 provides the numerical results. Finally, Section 7 concludes with a summary and discussion. Unless otherwise specified, we adopt geometric units, i.e., $c=G=1$, where $c$ is the vacuum speed of light and $G$ is the gravitational constant.
@@ 2 Review of Black Holes in Kalb-Ramond Gravity
$\omega=\frac{d\phi}{dt}=-\frac{g_{t\phi}}{g_{\phi\phi}}=\frac{2Mar-\frac{\varsigma a}{r^{-2(\varpi-1)/\varpi}}}{(r^{2}+a^{2})^{2}-a^{2}\sin^{2}\theta\Delta}.$ (6)
@@ 3.2 Synchrotron Radiation
$e_{(k)}^{\mu}=-\left(\frac{k^{\mu}}{u^{\nu}k_{\nu}}+u^{\mu}\right),\quad e_{(b)}^{\mu}=\frac{b^{\mu}}{b},$ (29)
@@ 4 Accretion Disk Models
For the non-polarized case, two geometrically thick and optically thin accretion disk models can be considered: the RIAF model [67] and the BAAF model [68, 69]. The former is highly consistent with general relativistic magnetohydrodynamic (GRMHD) simulations and has successfully reproduced the overall morphological features of M87* [9]. However, it neglects outflows, non-thermal particles, and the full GRMHD dynamics, thus limiting its application in polarization studies. The latter assumes that the fluid near the event horizon is primarily accelerated by gravity and provides explicit expressions for the thermodynamic variables and magnetic field configuration, offering a better description of the morphology and dynamics of the geometrically thick accretion flow in the near-horizon region of the black hole.
4.1 RIAF Model
@@
where the dimensionless quantity $\rho=n_{e}(m_{p}c^{2})$ represents the fluid mass density. For the accretion disk model considered in this paper, the order of magnitude of $\sigma$ is $\sigma\sim 0.1$ [70]. For the RIAF model, both isotropic radiation (35) and anisotropic radiation (33) can be considered.
@@
4.2 BAAF Disk Model
The BAAF model is a steady-state, axisymmetric accretion model [69, 68]. In this model, the fluid is treated as electrically neutral and the plasma is assumed to be fully ionized into electrons and protons. The accreting particle is confined to constant $\theta$ surface, i.e., $u^{\theta}\equiv 0$. In this case, the mass conservation equation is
@@ 4.2 BAAF Disk Model
$\Omega_{b}$ is the angular velocity of the magnetic field, and in this paper, we take $\Omega_{b}=0.3\Omega_{h}$, where $\Omega_{h}={a}/(2r_{h})$ is the angular velocity of the rotating black hole. For the BAAF model, only anisotropic radiation (33) is considered. For the fluid’s four-velocity, we can use the ballistic approximation, where the fluid moves along the geodesics [69], or the infalling motion mentioned in (47). We will choose the later in this paper.
@@ 6 Numerical Results
6.1 RIAF Model under Isotropic Radiation
We first discuss the impact of the spin parameter $a$ on the RIAF model under isotropic radiation. We set the spontaneous Lorentz violating parameters $\varpi=\varsigma=0.1$, the observer inclination angle is $\theta_{o}=65^{\circ}$. The numerical results are shown in Figure 1. We observe a bright ring structure in all the images, corresponding to higher-order images, where photons orbit the black hole one or more times before reaching the observer. These higher-order images are a result of the strong gravitational lensing effect, where the black hole’s gravity bends the light path, allowing some photons to bypass the black hole and reach the observer. Beyond this ring structure, the region with nonzero intensity corresponds to the primary image, where photons travel directly from the accretion disk to the observer. Notably, for all parameters, there is a region of decreasing intensity inside the higher-order images, which originates from the black hole’s event horizon. In the case of a geometrically thin accretion disk, the particle is confined to the equatorial plane, so the event horizon appears as a clearly defined black region in the image (the ”inner shadow”), which may be captured by the EHT [79]. However, for a geometrically thick accretion disk, this region may be obscured by radiation from outside the equatorial plane, making it difficult to distinguish. Compared to thin disks, thick disks are more physically realistic, suggesting that direct imaging of the black hole’s event horizon remains challenging.
@@ 6.1 RIAF Model under Isotropic Radiation
Figure 1: The impact of the spin parameter $a$ on the RIAF model under isotropic radiation. Fixed parameters are $\varpi=\varsigma=0.1$, and $\theta_{o}=65^{\circ}$.
@@
Figure 3: Black hole shadow images under the RIAF model with isotropic radiation. The parameters are fixed at $a=0.1$, $\theta_{o}=65^{\circ}$.
@@
Figure 4: Intensity distributions along the x-axis and y-axis under the RIAF model with isotropic radiation. The red, green, blue, and orange colors correspond to $\varsigma=0.69,0.79,0.89,0.99$, respectively. The parameters are fixed at $a=0.1$, $\varpi=0.9$, $\theta=65^{\circ}$.
@@
Figure 5: Intensity distributions along the x-axis and y-axis under the RIAF model with isotropic radiation. The red, green, and blue colors correspond to $\varpi=0.3,0.6,0.9$, respectively. The parameters are fixed at $a=0.1$, $\varsigma=0.99$, $\theta_{o}=65^{\circ}$.
6.2 RIAF Model under Anisotropic Radiation
Figures 6 and 7 show the impact of the spin parameter $a$ on the RIAF model under anisotropic radiation, along with the corresponding blurred images. For isotropic radiation, when $a$ is small, the higher-order images are approximately circular, but under anisotropic radiation, they significantly deform into elliptical shapes. The main reason for this difference is the different fitting functions for $\mathcal{F}(x)$. Similar to Figure 1, as $a$ increases, the intensity on the left side of the ellipse increases. Figure 8 shows the impact of $\varpi$ and $\varsigma$ on the shadow image. It can be observed that increasing $\varpi$ and $\varsigma$ reduces the size of the ellipse. Figures 9 and 10 show the corresponding intensity distributions. For both the x-axis and the y-axis, under anisotropic radiation, the distance between the two peaks is smaller than in the case of isotropic radiation. The above analysis indicates that the radiation type affects both the shape and size of the higher-order images.
@@ 6.2 RIAF Model under Anisotropic Radiation
Figure 6: The impact of the spin parameter $a$ on the RIAF model under anisotropic radiation. The fixed parameters are $\varpi=\varsigma=0.1$, and $\theta_{o}=65^{\circ}$.
@@
Figure 8: Black hole shadow images under the RIAF model with anisotropic radiation. The fixed parameters are $a=0.1$, $\theta_{o}=65^{\circ}$.
@@
Figure 9: Intensity distributions along the x-axis and y-axis under the RIAF model with anisotropic radiation. The red, green, blue, and orange colors correspond to $\varsigma=0.69,0.79,0.89,0.99$, respectively. The fixed parameters are $a=0.1$, $\varpi=0.9$, $\theta=65^{\circ}$.
@@
Figure 10: Intensity distributions along the x-axis and y-axis under the RIAF model with anisotropic radiation. The red, green, and blue colors correspond to $\varpi=0.3,0.6,0.9$, respectively. The fixed parameters are $a=0.1$, $\varsigma=0.99$, $\theta_{o}=65^{\circ}$.
6.3 BAAF Model
Figure 11 shows the impact of the spin parameter $a$ on the BAAF model, and Figure 12 presents the corresponding blurred images. When $a$ is small, the higher-order images appear circular, similar to the case of isotropic radiation. However, unlike the isotropic case, the higher-order images in the BAAF model feature a complete region of intensity reduction inside them. This indicates that, compared to the RIAF model, the radiation from outside the equatorial plane in the BAAF model has a weaker effect in obscuring the event horizon. Comparing the RIAF model and the BAAF model, we find that the latter is closer to the EHT observational results. As shown in Figure 13, the impact of $\varpi$ and $\varsigma$ on the higher-order images in the BAAF model is similar to that in the RIAF model. However, the region corresponding to the primary image is significantly smaller in the BAAF model. This conclusion is also evident from the intensity distribution diagrams in Figures 14 and 15. For both the x-axis and y-axis, the peaks corresponding to the higher-order images in the BAAF model are sharper, and the intensity rapidly approaches zero beyond the peaks. This reflects that, in the BAAF model, the intensity contribution primarily originates from the higher-order images. In contrast, in the RIAF model, the intensity of the primary image outside the higher-order images remains relatively high.
@@ 6.3 BAAF Model
Figure 11: The impact of the spin parameter $a$ on the BAAF model. The fixed parameters are $\varpi=\varsigma=0.1$, and $\theta_{o}=65^{\circ}$.
@@
Figure 13: Black hole shadow image under the BAAF model. The parameters are fixed at $a=0.1,\theta_{o}=65^{\circ}$.
@@
Figure 14: Intensity distribution along the x-axis and y-axis for the BAAF model. The red, green, blue, and orange colors correspond to $\varsigma=0.69,0.79,0.89,0.99$, respectively. The parameters are fixed as $a=0.1,\varpi=0.9,\theta=65^{\circ}$.
@@
Figure 15: Intensity distribution along the x-axis and y-axis for the BAAF model. The red, green, and blue colors correspond to $\varpi=0.3,0.6,0.9$, respectively. The parameters are fixed as $a=0.1,\varsigma=0.99,\theta_{o}=65^{\circ}$.
6.4 Polarization Images Illuminated by BAAF Disk
For the Polarization imaging, we will take the BAAF model as an example. Figure 16 shows the numerical results for the Stokes parameters $\tilde{I}_{o}$, $\tilde{Q}_{o}$, $\tilde{U}_{o}$, and $\tilde{V}_{o}$. Here, $\tilde{I}_{o}$ reflects the intensity distribution, with arrows representing the linear polarization vector $\vec{f}$. The color and direction of the arrows correspond to the degree of linear polarization $\mathcal{P}_{o}$ and the electric vector position angle $\Phi_{\mathrm{EVPA}}$. Since $\vec{f}\perp\vec{b}$, it can be inferred that the magnetic field is roughly radially distributed. The parameters $\tilde{Q}_{o}$ and $\tilde{U}_{o}$ reach their peaks near the higher-order images and decay rapidly away from this region. The $\tilde{V}_{o}$ parameter indicates right-handed circular polarization on both sides of the higher-order images, with left-handed circular polarization in the remaining regions.